OpenStudy (australia10):

0.7141414 repeating (14) decimal to a fraction ? are both methods valid? 1000x=714.1414 10x=7.14 990x=707 707/990 Or 100x=71.41414 x = .71414 99x=70.7 70.7/99 x 10 both sides 707/990

OpenStudy (mathmate):

Yes, both are valid, because they basically do the same thing. The first one works with numbers 10 times bigger than the second. For a quick check, subtract 7 from 714 and divide by 990 (99 from 14 repeating, 0 from 7) to get 707/990. Example, 0.7423232323...=(7423-74)/9900=7349/9900=0.74232323....